# DGE-YOLO: Dual-Branch Gathering and Attention for Accurate UAV Object Detection

### v1

@@ Abstract
The rapid proliferation of unmanned aerial vehicles (UAVs) has highlighted the importance of robust and efficient object detection in diverse aerial scenarios. Detecting small objects under complex conditions, however, remains a significant challenge. Existing approaches often prioritize inference speed, leading to degraded performance when handling multi-modal inputs. To address this, we present DGE-YOLO, an enhanced YOLO-based detection framework designed to effectively fuse multi-modal information. Specifically, we introduce a dual-branch architecture for modality-specific feature extraction, enabling the model to process both infrared and visible images. To further enrich semantic representation, we propose an Efficient Multi-scale Attention (EMA) mechanism that enhances feature learning across spatial scales. Additionally, we replace the conventional neck with a Gather-and-Distribute module to mitigate information loss during feature aggregation. Extensive experiments on the Drone Vehicle dataset demonstrate that DGE-YOLO achieves superior performance over state-of-the-art methods, validating its effectiveness in multi-modal UAV object detection tasks.
Index Terms:
@@
self-attention, object detection, multimodal, UAV, YOLO
I Introduction
Unmanned aerial vehicle (UAV) object detection has become a critical component in aerial perception systems, enabling intelligent applications across surveillance, traffic monitoring, and disaster response. With increasing demand for real-time and high-accuracy detection in complex aerial scenarios, challenges such as small object localization, modality imbalance, and computational constraints become prominent. In particular, single-modality detection systems—based solely on either infrared or visible images—often suffer from incomplete feature representation under diverse environmental conditions.
@@ I Introduction
Recent advances in image generation and multi-modal learning have demonstrated the benefits of modality-specific modeling and information consistency [29, 28, 31]. Inspired by these insights, we explore how the complementary advantages of infrared and visible images can be harnessed to improve UAV-based object detection. Instead of treating the fusion of modalities as a preprocessing step, we propose a novel lightweight multi-modal detection framework, termed DGE-YOLO, that performs joint feature extraction, attention-guided fusion, and global context aggregation in an end-to-end manner.
Figure 1: Overview of the framework of three detection strategies. (a) Is a single-stage detection model with only single-mode input, (b) is a two-stage detection model connecting the upstream image fusion task and the downstream detection task. (c) It is a Feature level fusion dge-yolo
Existing object detectors can be broadly categorized into two groups. The first includes general-purpose architectures such as R-CNN [1], Fast R-CNN [2], Faster R-CNN [3], SSD [4], YOLO, and DETR [5]. While these methods offer a balance between speed and accuracy, they typically process single-modality inputs and are not tailored for UAV-based multi-modal tasks. The second group focuses on multi-modal fusion, where either early-stage image fusion or two-stage detection is employed. For example, Tar-DAL [6] first aligns images from different modalities and then applies a detection network, but such two-stage strategies often result in excessive computational overhead. More recent approaches like OAFA [7] introduce dedicated encoders for modality-specific and invariant feature extraction, showing promising results on UAV benchmarks such as DroneVehicle [8]. However, these models still struggle with efficient feature integration and effective information retention.
To address these issues, our proposed DGE-YOLO introduces three core improvements. First, we design a dual-branch backbone that processes infrared and visible inputs separately and fuses them across multiple semantic scales. This design draws upon the hierarchical conditioning principles explored in prior conditional generation frameworks [30, 29]. Second, we incorporate an Efficient Multi-scale Attention (EMA) module [9] to dynamically model long-range dependencies and enhance cross-scale feature representation. Third, we integrate a Gather-and-Distribute (GAD) mechanism [10] as the neck module, allowing global information to be aggregated and redistributed effectively across the feature hierarchy. Fig 1. illustrates the differences between traditional image-level fusion, existing two-stage models, and our proposed end-to-end framework. Extensive experiments on the DroneVehicle dataset show that DGE-YOLO achieves superior detection performance under various aerial scenarios, validating the benefits of lightweight multi-modal integration and attention-driven feature refinement. Our contributions can be summarized as follows:
  • $\bullet$
    We propose a dual-branch lightweight backbone for multi-modal UAV object detection, which processes infrared and visible streams independently before multi-level fusion, balancing accuracy and efficiency.
  • $\bullet$
    We introduce a multi-scale attention module (EMA) to enhance both local and global feature dependencies across spatial resolutions.
  • $\bullet$
    We adopt a Gather-and-Distribute mechanism as the neck to enable efficient cross-layer feature aggregation and semantic propagation.
II Related Work
II-A Multimodal Object Detection
Multimodal object detection integrates complementary signals from various sources, such as RGB, infrared, depth, or LiDAR, to improve robustness under challenging conditions. In autonomous driving, M3D-RPN [16] and DeepLiDAR [17] incorporate depth cues to enhance spatial reasoning. In UAV-based detection, OAFA [7] introduces modality-specific and invariant feature extraction, and achieves notable performance on the DroneVehicle dataset [8]. Beyond spatial sensors, cross-modal learning has expanded to richer modalities. Works like BLIP-2 [11], and LLaVA [13] fuse vision and language for generalized understanding, while PaLM-E [15] enables embodied reasoning across images, actions, and language. However, the large scale of such models limits their deployment in real-time aerial systems.
II-B Efficient Attention Mechanisms
Self-attention, initially introduced in natural language processing, has been widely adopted in vision models. Vision Transformer (ViT) [22] treats image patches as tokens, while DETR [5] formulates detection as a set prediction problem using global self-attention. Swin Transformer [23] mitigates the computational burden via windowed attention. Despite their success, full self-attention remains resource-intensive. Lightweight adaptations have emerged in other domains, such as progressive diffusion [37] and controllable generation [36], where localized attention is used to balance precision and efficiency. These efforts reflect the growing trend toward task-specific, computationally-aware attention mechanisms.
II-C Multi-Scale Feature Fusion
Multi-scale representation plays a vital role in object detection. FPN [18] introduces top-down and lateral connections to aggregate features across scales, and RetinaNet [19] enhances it with focal loss to address class imbalance. PANet [21] and BiFPN [20] further improve the bidirectional flow and feature refinement. Nevertheless, traditional pyramid networks often suffer from inefficiencies and information loss during fusion. To address this, recent works in remote sensing and aerial perception [43, 44, 45] explore alternative fusion schemes that better preserve semantics across scales. These include frequency-aware modeling, spatial alignment, and optimized path aggregation strategies, reflecting the growing need for efficient multi-scale feature design in resource-constrained detection systems.
@@ II-C Multi-Scale Feature Fusion
Figure 2: The architecture of the proposed DGE-YOLO. Where $P_{a}$ and $P_{b}$ respectively represent Feature extracted from infrared and visible images by backbone, while $P_{c}$ represents Feature after fusion. $P_{c}$ is divided into $B2$, $B3$, $B4$ and $B5$ for subsequent processing. Avgpool, Billinear and concat in Low-Gather-and-Distribute form Low-FAM, and the rest is Low-IFM. Similarly, avgpool and concat in High-Gather-and-Distribute form High-FAM, and the rest is High-IFM. Inject:Information Injection Module.
III Method
III-A Overview
@@
To address the challenges faced by drones in performing target detection tasks in complex scenarios, such as slow detection speed and high missed detection rates, this paper modifies the backbone network of the YOLOv8n model by implementing a dual-channel input and introducing the EMA self-attention mechanism. Additionally, the neck network structure is replaced with the Gather-and-Distribute mechanism to further enhance the model’s performance.
III-B Dual-Branch Feature Extraction Network
@@
Feature fusion based on multimodal images generally divided into three strategies: early fusion, middle fusion, and late fusion. Early fusion combines the feature vectors of each modality during the input phase, either by concatenating them or aggregating them globally into a unified input vector, effectively merging different modality data at the data layer. This method is simple and efficient for handling low-level features; however, it struggles to address the differences between modalities, leading to potential information loss and poor handling of multi-scale information, resulting in suboptimal feature fusion performance. Late fusion, on the other hand, processes the input information of the two modalities independently before merging the data. This approach allows for independent optimization of each branch but fails to capture the correlations between features at different levels, thus limiting its performance potential. Middle fusion involves feature fusion at specific feature layers, which avoids the shortcomings of early fusion, such as the inability to effectively handle high-level features, while preserving the independent information of each modality. It also captures the relationships between different layers more effectively. Compared to the original model, which outputs from three feature extraction layers, we concatenate the features from all four feature extraction layers, ensuring that the network can fully utilize features from different layers, thereby enhancing the model’s ability to handle multi-scale tasks.
III-C Effiicient Multi-Scale Attention Module
@@
Effiicient multi-Scale attention module avoids deeper sequential processing and large depth by utilizing parallel substructures. Initially, the input feature map $X$ is divided into G sub-feature groups, each of which learns different semantics. This feature grouping method not only strengthens the feature learning of semantic regions but also enables the model to distribute processing across multiple GPUs, while compressing noise in the process EMA uses three parallel paths to extract attention weight descriptors for the grouped feature maps. Two of the paths are $1\times 1$ branches, which perform one-dimensional global average pooling operations to encode features in two spatial directions. The third path is a $3\times 3$ branch, which employs a $3\times 3$ convolution to capture multi-scale feature representations. Through the parallel sub-networks, EMA can adjust the weights of different channels by encoding cross-channel information and precisely retain spatial structural information within the channels. EMA also enriches feature aggregation by providing a cross-space information aggregation method in different spatial dimensions. The output from the $1\times 1$ branch is encoded as global spatial information through two-dimensional global average pooling, while the output from the $3\times 3$ branch is directly transformed into the corresponding dimensional shape. These outputs are then aggregated via matrix dot product operations to generate a spatial attention map. Finally, the attention weights generated by the Sigmoid function are applied to aggregate the output feature maps within each group, capturing pixel-level pairwise relationships and emphasizing the global context of all pixels. The final output maintains the same size as the input, allowing it to be effectively stacked within the network structure. The overall architecture is shown in Fig.3.
Figure 3: EMA Module Network Architecture Diagram
III-D Gather-and-Distribute mechanism
The Gather-and-Distribute mechanism utilizes a unified feature alignment module (FAM) to collect information from all levels and then uses an information fusion module (IFM) to merge the aligned features while retaining global information. The generated information is then distributed back to each level using the Inject module, where a simple attention operation is employed to inject this information, thereby enhancing the detection capability of the branches and significantly improving the information fusion ability in the neck section.To enhance the model’s ability to handle multi-scale objects, the mechanism uses two branches: the low-stage gather-and-distribute (Low-GD) branch and the high-stage gather-and-distribute (High-GD) branch. These branches are designed to extract and fuse feature maps for large and small object sizes, respectively.processing through the Transformer module.
Stage And Distribute Branch. The Low-GD branch collects feature vectors obtained from concatenating the backbone network denoted as $B2$, $B3$, $B4$ and $B5$. These four features are fused together to better preserve high-resolution features, which are crucial for small object detection. The High-GD branch fuses features generated by the Low-GD branch, denoted as $P3$, $P4$, and $P5$, to retain the semantic information contained in the features.
[…]
Stage Feature Fusion Module. The low-stage information fusion module (Low-IFM) includes multi-layer reparameterized convolution blocks and segmentation operations. The results obtained from Low-FAM undergo a RepBlock operation to generate $F_{low\_fuse}$, which is then split along the channel dimension using the segmentation operation into $F_{i\_P3}$ and $F_{i\_P4}$. These are then fused with features from different layers. Similarly, the High-IFM (High-Stage Information Fusion Module) contains Transformer blocks and segmentation operations. The results from High-IFM are passed through the Transformer block to generate $F_{high\_fuse}$, which is then reduced in the number of channels via convolution to sum($C_{P4}$, $C_{P5}$). Finally, the channels are split into $F_{i\_N4}$ and $F_{i\_N5}$, which are fused with the current layer’s features for integration. These feature alignment and fusion mechanisms effectively ensure the preservation of multi-scale information while minimizing computational overhead, leading to better performance in detecting objects at various scales. $F$ represent features.
$F_{low\_fuse}=RepBlock(Low-FAM([B2,B3,B4,B5])).$ (1)
$F_{i\_P3},F_{i\_P4}=Split(F_{low\_fuse}).$ (2)
$F_{high\_fuse}=transformer(High-FAM([P3,P4,P5])).$ (3)
$F_{i\_N4},F_{i\_N5}=Split(F_{high\_fuse}).$ (4)
[…]
$F_{g\_act}=resize(Sigmoid(Conv_{act}(F_{i})).$ (5)
$F_{g\_embed}=resize(Conv_{g\_embed}(F_{i})).$ (6)
$F_{att_{fuse}}=Conv_{l\_embed}(F_{l})*F_{i\_act}+F_{g\_embed}.$ (7)
@@ III-D Gather-and-Distribute mechanism
TABLE I: Comparison with State-of-the-art Methods.THE RED, GREEN AND BLUE ROWS REPRESENT THE 1ST, 2ND AND 3RD PLACES.RESPECTIVELY.
@@
YOLOv8s[27] 11.1 28.4 96.1 74.6 62.1 95.6 63.9 78.4 56.1
YOLOv9s[32] 7.3 26.7 96.1 76.0 64.4 96.2 64.8 79.5 57.6
YOLOv10s[33] 8.0 24.5 96.0 74.8 61.6 95.9 64.8 78.6 56.5
YOLOv11s[34] 9.4 21.3 96.0 75.7 60.4 96.3 64.4 78.5 56.5
YOLOv12s[35] 9.2 21.2 96.0 76.0 63.9 95.9 63.4 79.0 56.7
@@
Halfway Fusion[39] - - 90.1 62.3 58.5 89.1 49.8 70.0 -
CIAN[40] - - 90.1 63.8 60.7 89.1 50.3 70.8 -
@@
MBNet[42] - - 90.1 64.4 62.4 88.8 53.6 71.9 -
TSFADet[46] - - 89.9 67.9 63.7 89.8 54.0 73.1 -
$c^{2}$ormer[47] - - 90.2 68.3 64.4 89.8 58.5 74.2 -
SLBAF-Net[48] - - 90.2 72.0 68.6 89.9 59.9 76.1 -
@@
Despite these advances, few studies can meet the challenge of UAV target detection under the condition of ignoring Feature information in different modes. In this paper, we propose a novel method DGE-YOLO, which can not only better deal with Feature in different modes, but also modify the Feature processing process to better integrate Feature information to achieve better model performance.
IV Experiment and Analysis
IV-A Datasets
Drone Vehicle. This paper evaluates our method on the Drone Vehicle dataset, which contains 28,439 pairs of RGB-Infrared images, covering various scenarios such as urban roads, residential areas, parking lots, and scenes from day to night. The dataset includes five object categories: car, truck, bus, freight car, and van, although it still exhibits some weak misalignment issues. Following the training approach in the original paper, we used 17,990 image pairs for training and 1,469 image pairs for validation. The evaluation metrics selected are mAP50 and mAP50:95.
IV-B Implementation Details
The experiments are carried out on a single NVIDIA RTX 4090 GPU with 24 GB of memory. We implement our algorithm with Pytorch toolbox and auto optimizer with a momentum of 0.937 and a weight decay of 0.0005. The initial learning rate is set to 0.01. The input size of our model is set to 640×640 in the preprocessing stage and the batch size is 16. The training epoch is set to 300. Since infrared images contain more complete target characteristics, this paper used the labels from the infrared images as the training and validation labels.
IV-C Evaluation Metrics
To thoroughly demonstrate the advanced nature of the model, the experimental section compares the mAP50 for each class as well as the overall mAP50 and mAP50:95 as evaluation metrics. Since some two-stage models and multimodal fusion models use different evaluation standards, this paper compares the mAP50:95 metric only with YOLO-related models. Compared to two-stage and multimodal fusion models, one-stage and single-modality models generally exhibit much faster detection speeds. Therefore, this paper also compares the model’s parameter count and FLOPs only with YOLO series models.
IV-D Comparison with State-of-the-art Methods
This paper compares our method with state-of-the-art single-modal and multimodal object detection models. The single-modal models include one-stage models such as RetinaNet, $S^{2}$et, YOLOv8-YOLOv12, and two-stage models like Faster R-CNN and RoITransformer. For the single-modal object detection models, we trained them using RGB and IR images from the dataset separately and provided the experimental results. Additionally, our model is compared with other multimodal models, including UA-CMDet, Halfway Fusion, CIAN, MBNet, AR-CNN, TSFADet, $c^{2}$ormer, SLBAF-Net and OAFA. To ensure fairness, the original experimental settings were adopted for each comparison method. The experimental results are shown in Table1. Among the single-modal detection models, the detection performance under the infrared modality is significantly higher than that under the visible light modality. Given that the newer versions of YOLO have demonstrated excellent performance in detection tasks, including drone detection, we selected the YOLOv8 and subsequent versions to validate the effectiveness of our model. Compared to previous detection models, our model exhibits a significant improvement in accuracy. Since the parameters and flops of the single-stage detection model are much smaller than those of the multimodal fusion model, no comparison is made in the experiment. Our model has shown substantial improvements in categories with previously lower accuracy, achieving the highest mAP50, and a significant boost in the mAP50:95 metric.The comparison results shown in TABLEI demonstrate that the proposed DGE-YOLO method outperforms many state-of-the-art single-modal and multimodal object detection methods.

### v2

# DGE-YOLO: Dual-Branch Gathering and Effiicient Attention for Accurate UAV Object Detection

@@ Abstract
The rapid proliferation of unmanned aerial vehicles (UAVs) has highlighted the importance of robust and efficient object detection in diverse aerial scenarios. Detecting small objects under complex conditions, however, remains a significant challenge.To address this, we present DGE-YOLO, an enhanced YOLO-based detection framework designed to effectively fuse multi-modal information. We introduce a dual-branch architecture for modality-specific feature extraction, enabling the model to process both infrared and visible images. To further enrich semantic representation, we propose an Efficient Multi-scale Attention (EMA) mechanism that enhances feature learning across spatial scales. Additionally, we replace the conventional neck with a Gather-and-Distribute(GD) module to mitigate information loss during feature aggregation. Extensive experiments on the Drone Vehicle dataset demonstrate that DGE-YOLO achieves superior performance over state-of-the-art methods, validating its effectiveness in multi-modal UAV object detection tasks.
@@
Index Terms—  self-attention, object detection, multimodal, UAV, YOLO
1 Introduction
@@
Unmanned aerial vehicle object detection [tang2024fourier, liu2024feature] has become a critical component in aerial perception systems. With increasing demand for real-time and high-accuracy detection in complex aerial scenarios, challenges such as small object localization, modality imbalance, and computational constraints become prominent. Recent advances in image generation and multi-modal learning have demonstrated the benefits of modality-specific modeling and information consistency [weng2024enhancing, li2024lr, qiao2022novel]. Inspired by these insights, we explore how the complementary advantages of infrared and visible images can be harnessed to improve UAV-based object detection. Instead of treating the fusion of modalities as a preprocessing step, we propose a novel lightweight multi-modal detection framework, termed DGE-YOLO, that performs joint feature extraction, attention-guided fusion, and global context aggregation in an end-to-end manner.
Fig. 1: Schematic diagrams of three types of detection frames.
@@
Fig. 2: The architecture of the proposed DGE-YOLO. Where $P_{a}$ and $P_{b}$ respectively represent Feature extracted from infrared and visible images by backbone, while $P_{c}$ represents Feature after fusion. $P_{c}$ is divided into $B2$, $B3$, $B4$ and $B5$ for subsequent processing. Avgpool, Billinear and concat in Low-Gather-and-Distribute form Low-FAM, and the rest is Low-IFM. Similarly, avgpool and concat in High-Gather-and-Distribute form High-FAM, and the rest is High-IFM.
To address these issues, our proposed DGE-YOLO introduces three core improvements. First, we design a dual-branch backbone that processes infrared and visible inputs separately and fuses them across multiple semantic scales. This design draws upon the hierarchical conditioning principles explored in prior conditional generation frameworks [shen2025imaggarment]. Second, we incorporate an EMA module [2023Efficient] to dynamically model long-range dependencies and enhance cross-scale feature representation. Third, we integrate a GD mechanism [NEURIPS2023_a0673542] as the neck module, allowing global information to be aggregated and redistributed effectively across the feature hierarchy. Fig 1. illustrates the differences between traditional image-level fusion, existing two-stage models, and our proposed end-to-end framework. Extensive experiments on the DroneVehicle [9759286] dataset show that DGE-YOLO achieves superior detection performance under various aerial scenarios. Our contributions can be summarized as follows: We propose a dual-branch lightweight backbone for multi-modal UAV object detection, which processes infrared and visible streams independently before multi-level fusion, balancing accuracy and efficiency. We introduce a EMA to enhance both local and global feature dependencies across spatial resolutions. We adopt a GD mechanism as the neck to enable efficient cross-layer feature aggregation and semantic propagation.
2 METHOD
2.1 Overview
@@
To address the challenges faced by drones in performing target detection tasks in complex scenarios, such as slow detection speed and high missed detection rates, this paper modifies the backbone network of the YOLOv8n model by implementing a dual-channel input and introducing the EMA self-attention mechanism. Additionally, the neck network structure is replaced with the GD mechanism to further enhance the model’s performance. As shown in Fig.2
2.2 Dual-Branch Feature Extraction Network
@@
Feature fusion based on multimodal images generally divided into three strategies: early fusion, middle fusion, and late fusion. Middle fusion involves feature fusion at specific feature layers, which avoids the shortcomings of early fusion, such as the inability to effectively handle high-level features, while preserving the independent information of each modality. It also addresses the issue of late fusion not being able to effectively capture the relationships between different layers. Compared to the original model, which outputs from three feature extraction layers, we concatenate the features from all four feature extraction layers, ensuring that the network can fully utilize features from different layers, thereby enhancing the model’s ability to handle multi-scale tasks.
2.3 Effiicient Multi-Scale Attention Module
Fig. 3: EMA Module Network Architecture Diagram.
@@ 2.3 Effiicient Multi-Scale Attention Module
Effiicient multi-Scale attention module avoids deeper sequential processing and large depth by utilizing parallel substructures. While compressing noise in the process EMA uses three parallel paths to extract attention weight descriptors for the grouped feature maps. The output from the $1\times 1$ branch is encoded as global spatial information through two-dimensional global average pooling, while the output from the $3\times 3$ branch is directly transformed into the corresponding dimensional shape. These outputs are then aggregated via matrix dot product operations to generate a spatial attention map. Finally, the attention weights generated by the Sigmoid function are applied to aggregate the output feature maps within each group, capturing pixel-level pairwise relationships and emphasizing the global context of all pixels. The final output maintains the same size as the input, allowing it to be effectively stacked within the network structure. Through the parallel sub-networks, EMA can adjust the weights of different channels by encoding cross-channel information and precisely retain spatial structural information within the channels. EMA also enriches feature aggregation by providing a cross-space information aggregation method in different spatial dimensions.The overall architecture is shown in Fig.3.
2.4 Gather-and-Distribute Mechanism
@@
Table 1: Comparison with State-of-the-art Methods.THE RED, GREEN AND BLUE ROWS REPRESENT THE 1ST, 2ND AND 3RD PLACES.RESPECTIVELY.
@@
YOLOv8s[yolov8_ultralytics] 11.1 28.4 96.1 74.6 62.1 95.6 63.9 78.4 56.1
YOLOv9s[wang2024yolov9] 7.3 26.7 96.1 76.0 64.4 96.2 64.8 79.5 57.6
YOLOv10s[THU-MIGyolov10] 8.0 24.5 96.0 74.8 61.6 95.9 64.8 78.6 56.5
YOLOv11s[yolo11_ultralytics] 9.4 21.3 96.0 75.7 60.4 96.3 64.4 78.5 56.5
YOLOv12s[yolo12] 9.2 21.2 96.0 76.0 63.9 95.9 63.4 79.0 56.7
@@
Halfway Fusion[BMVC2016_73] - - 90.1 62.3 58.5 89.1 49.8 70.0 -
CIAN[ZHANG201920] - - 90.1 63.8 60.7 89.1 50.3 70.8 -
@@
MBNet[10.1007/978-3-030-58523-5_46] - - 90.1 64.4 62.4 88.8 53.6 71.9 -
TSFADet[10.1007/978-3-031-20077-9_30] - - 89.9 67.9 63.7 89.8 54.0 73.1 -
$c^{2}$Former[10472947] - - 90.2 68.3 64.4 89.8 58.5 74.2 -
SLBAF-Net[10.1007/s11042-023-15333-w] - - 90.2 72.0 68.6 89.9 59.9 76.1 -
@@
Stage Feature Alignment Module. In the low-stage feature alignment module (Low-FAM), input features are downsampled using average pooling, and the features are adjusted to the smallest feature size within the group ($R_{B4}$ = 1/4RR, $R$ represent size) to obtain aligned features. In the high-stage feature alignment module (High-FAM), average pooling is used to adjust the dimensions of the input features to a uniform size of $R_{B5}$ = 1/8RR. The feature alignment module ensures efficient aggregation of information while minimizing the computational complexity of subsequent.
Stage Feature Fusion Module. The low-stage information fusion module (Low-IFM) includes multi-layer reparameterized convolution blocks and segmentation operations. The results obtained from Low-FAM undergo a RepBlock operation to generate $F_{low\_fuse}$, which is then split along the channel dimension using the segmentation operation into $F_{i\_P3}$ and $F_{i\_P4}$. These are then fused with features from different layers. Similarly, the High-Stage Information Fusion Module(High-IFM) contains Transformer blocks and segmentation operations. The results from High-IFM are passed through the Transformer block to generate $F_{high\_fuse}$, which is then reduced in the number of channels via convolution to sum($C_{P4}$, $C_{P5}$). Finally, the channels are split into $F_{i\_N4}$ and $F_{i\_N5}$, which are fused with the current layer’s features for integration.
[…]
$F_{g\_act}=resize(Sigmoid(Conv_{act}(F_{i})).$ (1)
$F_{g\_embed}=resize(Conv_{g\_embed}(F_{i})).$ (2)
$F_{att_{fuse}}=Conv_{l\_embed}(F_{l})*F_{i\_act}+F_{g\_embed}.$ (3)
3 Experiment and Analysis